OpenStudy (anonymous):

A student scores 73 out of 100 on a test. If the maximum score on the next test is also 100 points, what score does the student need to maintain at least an average of 78?

OpenStudy (baldymcgee6):

(73 +x)/2 = 78

OpenStudy (baldymcgee6):

Solve for x :)

OpenStudy (anonymous):

\[\frac{ 73+x }{ 2} do you mean that?\]

OpenStudy (baldymcgee6):

\[\frac{ 73+x }{ 2 }= 78\] Yes, they are equivalent.

OpenStudy (anonymous):

\[\frac{ 73+x }{ 2 }=78\]\[73x+2=78\times2\]\[73+x=156\]\[x=83\]

OpenStudy (anonymous):

@EchaSetiaWijaya ahhh! that makes sense to me. i didn't understand it written in a fraction but I understand 73x + 2 = 78 x 2.

OpenStudy (anonymous):

@baldymcgee6 @EchaSetiaWijaya thanks for your help!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

sorry, for the second step this is the right on i mean 73 + x = 78 * 2